f(x)=2-6x,g(x)=-8+6
find (f+g)(x)

Answer by stanbon(75887)

f(x)=2-6x,g(x)=-8+6x
(f+g)(x) = f(x) + g(x) = 2-6x -8+6x = -6
=====================
Cheers,
Stan H.